# SoK: On Finding Common Ground in Loss Landscapes Using Deep Model Merging Techniques

### v1

@@ IX Conclusion & Future Work
We have surveyed a range of model merging techniques and extracted insights from their empirical observations to characterize several aspects of the underlying structure of loss landscapes—namely mode convexity, mode determinism, mode directedness, and mode connectivity. We discussed the intimate connections between these observations and model interpretability and security, offering analogs for traditional activation space analysis in weight space.
@@
Future work exploring the intersection of these fields might focus on understanding the relationship between the objective geometry of hierarchically related tasks, with the goal of understanding why semantically similar tasks produce solutions close together in parameter space, even when their raw data distributions (over, say, tokens) might seem quite dissimilar in principle [82]. Such investigations would touch on the ability of neural networks to abstract and generalize across tasks in a meaningful way. One could also attempt to “orthogonalize” task-specific directions in weight space to obtain interpretable decompositions over directions in parameter space as an alternative to activation-based methods. Other extensions might further explore methods for continual learning, few-shot learning, or parameter-efficient fine-tuning through model merging techniques.
@@
A concerning implication of our work is that, due to the principle of mode determinism, models fine-tuned from pre-trained checkpoints—particularly those fine-tuned on similar tasks—will have similar representations and lie close together in parameter space. Therefore, models stemming from freely available pre-trained checkpoints might be especially vulnerable to adversaries attempting to replicate model behavior [9, 147, 148]. This could lead to problems, given that the dominant paradigm for model deployment is currently predicated on extensive pre-training and domain-specific fine-tuning.
Acknowledgments
@@
This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112.

### v2

@@ 8. Conclusion & Future Work
We have surveyed a range of model merging works and extracted insights into the underlying structure of loss landscapes—namely by characterizing mode convexity, mode determinism, mode directedness, and mode connectivity. We discussed the intimate connections between these observations and model interpretability and robustness, offering analogs for traditional activation space analysis in weight space.
@@
Future work exploring the intersection of these fields might focus on understanding the relationship between the objective geometry of hierarchically related tasks to understand why semantically similar tasks produce solutions close together in parameter space, even when their raw data distributions (over, say, tokens) might seem quite dissimilar (44). Such investigations would touch on the ability of neural networks to abstract and generalize across tasks in a meaningful way. One could also attempt to “orthogonalize” task-specific directions in weight space to obtain interpretable decompositions over directions in parameter space as an alternative to activation-based methods such as Distributed Alignment Search (42), Sparse Autoencoders (13), and the Interaction Basis (15). Other extensions might further explore methods for continual learning, few-shot learning, or parameter-efficient fine-tuning through model merging techniques.
@@
A concerning implication of our work is that, due to the principle of mode determinism, models fine-tuned from pre-trained checkpoints—particularly those fine-tuned on similar tasks—will have similar representations and lie close together in parameter space. Therefore, models stemming from freely available pre-trained checkpoints might be especially vulnerable to adversaries attempting to replicate model behavior (105; 99; 10). This seems a significant risk, given that the dominant paradigm for model deployment is currently predicated on extensive pre-training and domain-specific fine-tuning. Further work exploring how to secure open-source models is paramount to ensuring a healthy open-source environment for deep learning models.
Acknowledgements.
@@
This material is based on work supported by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research, Department of Energy Computational Science Graduate Fellowship under Award Number DE-SC0023112 and by the U.S. Department of Energy, Office of Science, Advanced Scientific Computing Research, under Contract DE-AC02-06CH11357.